# Detection of Aerial Spoofing Attacks to LEO Satellite Systems via Deep Learning

### v1

@@ Abstract
Detecting spoofing attacks to Low-Earth-Orbit (LEO) satellite systems is a cornerstone to assessing the authenticity of the received information and guaranteeing robust service delivery in several application domains. The solutions available today for spoofing detection either rely on additional communication systems, receivers, and antennas, or require mobile deployments. Detection systems working at the Physical (PHY) layer of the satellite communication link also require time-consuming and energy-hungry training processes on all satellites of the constellation, and rely on the availability of spoofed data, which are often challenging to collect. Moreover, none of such contributions investigate the feasibility of aerial spoofing attacks launched via drones operating at various altitudes. In this paper, we propose a new spoofing detection technique for LEO satellite constellation systems, applying anomaly detection on the received PHY signal via autoencoders. We validate our solution through an extensive measurement campaign involving the deployment of an actual spoofer (Software-Defined Radio) installed on a drone and injecting rogue IRIDIUM messages while flying at different altitudes with various movement patterns. Our results demonstrate that the proposed technique can reliably detect LEO spoofing attacks launched at different altitudes, while state-of-the-art competing approaches simply fail. We also release the collected data as open source, fostering further research on satellite security.
Index Terms:
Wireless Security; Satellite Security; Artificial Intelligence for Security; Mobile Security; Drones.
I Introduction
Low-Earth Orbit (LEO) satellite communication systems have recently gained renewed momentum in Internet of Things (IoT) research and Industry, thanks to the enhanced simplicity and low cost in technology access [1]. Several commercial IoT-based applications are also benefiting from the integration of such a technology, with new satellite-based services emerging like real-time data collection, predictive maintenance, asset tracking, cold-chain management, and resource exploration, to name a few [2].
@@ I Introduction
Contribution. In this paper, we provide the following contributions: (i) we demonstrate the technical feasibility of aerial spoofing attacks to LEO satellite systems, where an adversary uses drones operating at various altitudes (beyond Line-of-Sight) to stealthily inject spoofed messages to a target satellite receiver; (ii) we demonstrate that existing solutions to detect LEO spoofing attacks fail to discover such attacks, misclassifying spoofed messages as legitimate ones; (iii) we design a new solution for aerial spoofing attacks to LEO satellite systems, using a sparse Autoencoder (AE) to reliably detect anomalies in the received Physical (PHY)-layer signals; and finally, (iv) through extensive real-world experimentation involving actual drones and Software-Defined Radios, we demonstrate the robustness and effectiveness of our proposed solution and its superior performances against the state of the art. To the best of our knowledge, this paper is the first to consider real-world aerial spoofing attacks to LEO satellite systems and to propose the usage of a sparse AE on a standalone receiver to detect such attacks. We also release the dataset collected for this work at [12], enabling further research on LEO satellite links modeling, monitoring, and security.
@@
  • While previous research used binary classification for spoofing detection, in this paper, we demonstrate the feasibility of detecting spoofing attacks to LEO satellite systems using one-class classification via autoencoders, thus requiring the availability of only regular communication samples at training time.
  • While previous research demonstrated the effectiveness of fading fingerprinting by only considering terrestrial attacks, in this paper, we consider a new attack strategy, i.e., aerial spoofing attacks launched through drones.
@@
Roadmap. This paper is organized as follows. Sec. II introduces the preliminaries, Sec. III reviews related work, Sec. IV introduces the system and adversary model, Sec. V provides the details of our solution, Sec. VI describes our data collection, Sec. VII provides the results, and finally, Sec. VIII draws the conclusion.
II Background
@@
This section introduces preliminary concepts for this manuscript, i.e., digital modulation principles, used as an input to detecting spoofing attacks, the IRIDIUM satellite infrastructure, used for our experimental testing phase, and AEs, used as part of our detection solution.
@@
Digital (De-)Modulation. Modern wireless communication technologies use digital modulation to make bit-strings suitable for transmission over the Radio Frequency (RF) link. Such signals are typically expressed in the form of a complex number $I+jQ$, namely In-Phase Quadrature (IQ) samples, being $I$ the In-Phase and $Q$ the Quadrature component. These components are (90 degrees) phase-shifted by the transmitter before transmission so that they do not interfere with each other while propagating over the wireless channel. The original IQ value (and bitstream) can be recovered by the receiver by multiplying the received signal via two orthogonal tones [14]. The number of possible symbols $s$ defined by the modulation scheme, i.e., $log_{2}{s}$ bits/sec, impacts the throughput of the communication link. In addition, $s$ defines the tolerance of the communication link to noise. In fact, the IQ symbols at the receiver are affected by noise $n$, introduced jointly by the electronics of the transmitter and the communication channel. The receiver simply maps the received IQ symbol $x=I+jQ+n$ to the closer expected one, i.e., the symbol characterized by the minimum error. The higher the noise power, the higher the probability that a symbol is distorted enough to get closer to another one, causing a decoding error. Thus, communication links affected by high noise, such as satellite links, use lower-order schemes, e.g., Binary Phase Shift Keying (BPSK) ($s=2$) and QPSK ($s=4$), trading off throughput with robustness to noise. Also, note that Commercial-Off-The-Shelfs cannot easily control the value of the IQ samples. Indeed, the IQ samples received by a device are affected by both manufacturing inaccuracies in the transmitter hardware, which are inherent to such devices, and fading phenomena caused by the wireless channel [15]. In our paper, we analyze the distortion affecting the IQ samples received by a target user to detect spoofing attacks (see Sec. V).
@@
IRIDIUM. IRIDIUM is a LEO satellite constellation operated by IRIDIUM Communications Inc, offering global satellite phone and data coverage currently to more than 2 millions users [16][17]. IRIDIUM enables users to communicate from anywhere on Earth, including remote areas where traditional terrestrial communication networks typically cannot provide connectivity. The constellation includes $66$ LEO satellites orbiting at $800$ km all over the Earth’s surface and moving from north to south [18]. The satellites transmit messages in the frequency band $[1,618.75-1,626.5]$ MHz using the modulation scheme Differential (DQPSK), which can be received through dedicated satellite receiving stations provided by authorized resellers such as Motorola and Kyocera. IRIDIUM features two main types of communication channels, i.e., system overhead and bearer service channels [19]. Some channels are authenticated, while others are not authenticated and not encrypted, allowing for spoofing attacks. Unauthenticated and unencrypted channels, e.g., IRIDIUM Ring Alert (IRA), are used to broadcast information about a specific satellite. The information delivered by satellites on such channels can be detected and decoded using a SDR, a generic active antenna, and a software development kit like gr-iridium [20]. IRIDIUM satellites serve primarily as a communication system. However, the network can also be used for positioning purposes, e.g., through systems like DDK Positioning [21]. IRIDIUM also features a dedicated IRIDIUM IoT service, specifically dedicated to Machine-to-Machine (M2M) applications like transportation, agriculture, oil and gas, utilities, and construction, among many others [22]. In this paper, we use IRIDIUM as the reference LEO satellite constellation to test spoofing attacks because of its popularity, number of users, and availability of an open-source reception module for SDRs to receive real messages, enabling our experimental data collection and analysis.
@@
Note that AEs can also be used as a one-class classification tool, i.e., they can be trained only on samples from one class and then used to detect anomalies as outliers. This is particularly useful when dealing with aerial spoofing attacks. Indeed, collecting attack samples (anomalies) is particularly challenging because of the equipment, technical difficulties required to assemble the experimental setup and launch the attack, and finally, law and safety regulations, which often prevent the collection of such data. Recent related work, such as the one by the authors in [26], also identified experimentally the advantage of autoencoder-based models when a limited number of anomalous samples is available. We use this feature as part of our solution (see Sec. V).
Paper Location Independence Single Receiver Support for Static Receivers No crypto No Multiple Sources Applied on Satellites No Binary Classification Tests with Aerial Adversary
@@
TABLE I: Related work on Spoofing Detection.
III Related Work
@@
Spoofing of satellite messages is usually associated with Global Navigation Satellite System (GNSS) spoofing, i.e., the falsification of positioning and navigation messages emitted by systems such as GPS, GALILEO, and GLONASS [35]. Although secure versions of such communication technologies exist and are used by the military, for backwards-compatibility reasons, civilian systems mostly use their former versions, which are insecure by design. In this context, He et al. in [36] also identified that, similarly to GNSS technologies like GPS, LEO satellite networks typically use a dedicated downlink pilot channel for broadcasting channel status, user management information, call information, etc., as exemplified by Iridium. Attackers can imitate the dedicated pilot channel to broadcast false information to legitimate users, causing network paralysis. However, they did not provide valid countermeasures for such attacks. One common strategy to detect spoofing attacks involves the exploration of alternative sources, such as military GNSS signals [30], terrestrial communication systems [37], or the integration of data from additional GNSS constellations, such as Iridium [18]. These methodologies utilize the correlation between the information collected in distinct channels to assess the authenticity of GNSS signals. However, they depend on the availability and authenticity of an alternative source of information. Some other proposals, such as [29], use multiple receivers to detect spoofing attacks, leveraging spatial and channel diversity and the use of Channel State Information (CSI)-based fingerprints of the received signal. This approach has also been used by the authors in [13] and [31] for WiFi signals, and is overall very promising for indoor spoofing detection. However, as shown in Sec. VII-B, it hardly applies to GNSS spoofing due to high channel noise. To create channel diversity without several receivers, the authors in [32] use a mobile Global Positioning System (GPS) receiver. However, such solutions require mobile receivers and usually incur a non-negligible delay, necessary to actually move from one location to another. Recent popular methods to detect (GNSS) spoofing attacks involve the use of Radio Frequency Fingerprinting (RFF), i.e., the creation of a transmitter fingerprint that helps distinguish legitimate from rogue ones. RFF has been applied successfully for terrestrial communication technologies [38], LEO [8] and GPS satellites [27]. However, it requires significant training time and does not scale up with the number of legitimate satellites, which is typically large for LEO constellations. Other proposals in the GNSS domain, such as the one by Borhani et al. in [39], use GNSS-specific metrics, such as Cross Ambiguity Function (CAF), to detect spoofing attacks. Being such metrics specific to GNSS technologies, they cannot be used for LEO satellite spoofing detection. Some more recent proposals, such as [9][10], and [28], use IQ samples to detect spoofing attacks affecting both the LEO and GNSS constellations. While the authors in [28] detect anomalies in the Quadrature component of IQ signals, the proposals in [9] and [10] apply binary classification to both I and Q components to detect spoofing and require knowledge of both spoofed and legitimate samples. Binary classification is also required by Kumar et al. in [11], which applied Long Short-Term Memory (LSTM) to noise data via simulations to detect aerial spoofing. It is also worth mentioning the contributions by Abdrabou et al. in [33] and  [34], using Doppler frequency shift (DS) and received power (RP) features to detect spoofing attacks to satellite systems. However, none of these solutions have been tested with actual spoofing attacks carried out by aerial attackers, such as drones.
@@
Thus, as summarized in Table I, the current literature misses actual real-world tests involving spoofing attacks from aerial adversaries. Moreover, existing methods cannot provide a scalable mechanism independent of the number of satellites that comprise the LEO constellation. In Sec. VII-B, we also compare our solution to the solutions proposed in [10] and [8], demonstrating superior performance.
IV System and Adversary Model
@@
Without loss of generality, we consider a satellite transmitter located in the LEO, emitting RF messages according to the IRIDIUM communication system, as discussed in Sec. II. Thus, legitimate signals emitted by the satellite follow the modulation scheme QPSK with differential encoding, a.k.a. DQPSK, and are transmitted in the frequency band $[1,618.725-1,626.5]$ MHz. The user features a receiving device able to receive and decode IRIDIUM satellite messages, e.g., for communication in hard-to-reach locations. We consider a receiver deployed outdoors for maximum satellite visibility. We do not make any additional assumptions on the deployment of the receiver, which can be either static or mobile. However, we consider that the user does not feature any additional Internet connection than the one possibly provided by the LEO satellite system. Thus, the user cannot resort to any other service than the IRIDIUM communication system to verify the legitimacy of the information provided through such a system. This consideration applies to many IRIDIUM deployments, as the system is typically used to enable communications in remote and hard-to-reach locations. Moreover, note that we require raw physical-layer IQ data to detect spoofing attacks. Such data are usually not available when using COTSs, while they are available by using SDRs, as we did in our experimental campaign. Thus, users willing to deploy our solution require a SDR to collect the data aside to the IRIDIUM receiver provided by the service provider.
@@
V Satellite Spoofing Detection
In this section, we describe the methodology used to detect spoofing attacks to LEO satellite communication systems. Our proposed methodology for spoofing detection is divided into two phases, i.e., image generation from physical layer information, i.e., IQ samples, (see Sec. V-A) and spoofing detection (Sec. V-B). Tab. II summarizes our main notation.

### v2

@@ Abstract
Detecting spoofing attacks to Low-Earth-Orbit (LEO) satellite systems is a cornerstone to assessing the authenticity of the received information and guaranteeing robust service delivery in several application domains. The solutions available today for spoofing detection either rely on additional communication systems, receivers, and antennas, or require mobile deployments. Detection systems working at the Physical (PHY) layer of the satellite communication link also require time-consuming and energy-hungry training processes on all satellites of the constellation, and rely on the availability of spoofed data, which are often challenging to collect. Moreover, none of such contributions investigate the feasibility of aerial spoofing attacks launched via drones operating at various altitudes. In this paper, we first show experimentally the viability and effectiveness of spoofing attacks to LEO satellite systems using aerial attackers deployed on drones. We also propose a new spoofing detection technique, relying on pre-processing raw physical-layer signals into images and then applying anomaly detection on such images via autoencoders. We validate our solution through an extensive measurement campaign involving the deployment of an actual spoofer (Software-Defined Radio) installed on a drone and injecting rogue IRIDIUM messages while flying at different altitudes with various movement patterns. Our results demonstrate that the proposed technique can reliably detect LEO spoofing attacks launched at different altitudes, while state-of-the-art competing approaches simply fail. We also release the collected data as open source, fostering further research on satellite security.
keywords:
Wireless Security , Satellite Security , Artificial Intelligence for Security , Mobile Security , Drones
journal: Computer Networkst1t1footnotetext: This work has been partially supported by the INTERSECT project, Grant No. NWA.1162.18.301, funded by the Netherlands Organization for Scientific Research (NWO). Any opinions, findings, conclusions, or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of NWO. Moreover, this publication was made possible by the NPRP12C-0814-190012-SP165 awards from the Qatar National Research Fund (a member of Qatar Foundation).
[…]
[1]organization=Eindhoven University of Technology (TU/e), city=Eindhoven, country=Netherlands
[…]
[2]organization=Hamad Bin Khalifa University (HBKU), College of Science and Engineering (CSE), city=Doha, country=Qatar
1 Introduction
Low-Earth Orbit (LEO) satellite communication systems have recently gained renewed momentum in Internet of Things (IoT) research and Industry, thanks to the enhanced efficiency, simplicity, and low cost in technology access [1] [2]. Several commercial IoT-based applications are also benefiting from the integration of such a technology, with new powerful satellite-based services emerging like real-time data collection, predictive maintenance, asset tracking, cold-chain management, and resource exploration, to name a few, and new communication paradigms appearing, e.g., Mobile Edge Computing [3].
@@ 1 Introduction
Contribution. In this paper, we provide the following contributions: (i) we demonstrate the technical feasibility of aerial spoofing attacks to LEO satellite systems, where an adversary uses drones operating at various altitudes (beyond Line-of-Sight) to stealthily inject spoofed messages to a target satellite receiver; (ii) we demonstrate that existing solutions to detect LEO spoofing attacks fail to discover such attacks, misclassifying spoofed messages as legitimate ones; (iii) we design a new solution for aerial spoofing attacks to LEO satellite systems, pre-processing raw IQ samples into images and using a sparse Autoencoder (AE) to reliably detect anomalies in the received signals; and finally, (iv) through extensive real-world experimentation involving actual drones and Software-Defined Radios, we demonstrate the robustness and effectiveness of our proposed solution and its superior performances against the state of the art. To the best of our knowledge, this paper is the first to consider real-world aerial spoofing attacks to LEO satellite systems and to propose an ad-hoc pre-processing technique of physical-layer information enabling attack detection via sparse AE. Although similar techniques have been applied for identifying specific satellites, e.g., in [9], we demonstrate experimentally that satellite link fingerprinting is better suited for detecting spoofing attacks compared to identifying single satellites. We also release the dataset collected for this work at [13], enabling further research on LEO satellite links modeling, monitoring, and security.
@@
  • 2.
    While previous research leveraged binary classification for spoofing detection, in this paper, we demonstrate the feasibility of detecting spoofing attacks to LEO satellite systems using a pre-processing technique transforming physical-layer information into images, enabling spoofing detection through one-class classification via autoencoders, thus requiring the availability of only regular communication samples at training time.
  • 3.
    While previous research demonstrated the effectiveness of fading fingerprinting by only considering terrestrial attacks, in this paper, we consider a novel attack strategy, i.e., aerial spoofing attacks launched through drones.
  • 4.
@@
Roadmap. This paper is organized as follows. Sec. 2 introduces the preliminaries, Sec. 3 reviews related work, Sec. 4 introduces the system and adversary model, Sec. 5 provides the details of our solution, Sec. 6 describes our data collection, Sec. 7 provides the results, Sec. 8 discusses implications and limitations of our research, and finally, Sec. 9 draws the conclusion.
2 Background
@@
This section introduces preliminary concepts, i.e., digital modulation principles, used as an input to detecting spoofing attacks, the IRIDIUM satellite infrastructure, used for our experimental testing phase, and AEs, used as part of our detection solution.
@@
Digital (De-)Modulation. Modern wireless communication technologies use digital modulation to make bit-strings suitable for transmission over the Radio Frequency (RF) link. Such signals are typically expressed in the form of a complex number $I+jQ$, namely In-Phase Quadrature (IQ) samples, being $I$ the In-Phase and $Q$ the Quadrature component. These components are (90 degrees) phase-shifted by the transmitter before transmission so that they do not interfere with each other while propagating over the wireless channel. The original IQ value (and bitstream) can be recovered by the receiver by multiplying the received signal via two orthogonal tones [14]. The number of possible symbols $s$ defined by the modulation scheme, i.e., $log_{2}{s}$ bits/sec, impacts the throughput of the communication link. In addition, $s$ defines the tolerance of the communication link to noise. In fact, the IQ symbols at the receiver are affected by noise $n$, introduced jointly by the electronics of the transmitter and the communication channel. The receiver simply maps the received IQ symbol $x=I+jQ+n$ to the closer expected one, i.e., the symbol characterized by the minimum error. The higher the noise power, the higher the probability that a symbol is distorted enough to get closer to another one, causing a decoding error. Thus, communication links affected by high noise, such as satellite links, use lower-order schemes, e.g., Binary Phase Shift Keying (BPSK) ($s=2$) and QPSK ($s=4$), trading off throughput with robustness to noise. Also, note that Commercial-Off-The-Shelfs cannot easily control the value of the IQ samples. Indeed, the IQ samples received by a device are affected by both manufacturing inaccuracies in the transmitter hardware, which are inherent to such devices, and fading phenomena caused by the wireless channel [15]. In our paper, we analyze the distortion affecting the IQ samples received by a target user to detect spoofing attacks (see Sec. 5).
@@
IRIDIUM. IRIDIUM is a LEO satellite constellation operated by IRIDIUM Communications Inc, offering global satellite phone and data coverage currently to more than 2 millions users [16], [17]. IRIDIUM enables users to communicate from anywhere on Earth, including remote areas where traditional terrestrial communication networks typically cannot provide connectivity. The constellation includes $66$ LEO satellites orbiting at $800$ km all over the Earth’s surface and moving from north to south [18]. The satellites transmit messages in the frequency band $[1,618.75-1,626.5]$ MHz using the modulation scheme Differential (DQPSK), which can be received through dedicated satellite receiving stations provided by authorized resellers such as Motorola and Kyocera. IRIDIUM features two main types of communication channels, i.e., system overhead and bearer service channels [19]. Some channels are authenticated, while others are not authenticated and not encrypted, allowing for spoofing attacks. Unauthenticated and unencrypted channels, e.g., IRIDIUM Ring Alert (IRA), are used to broadcast information about a specific satellite. The information delivered by satellites on such channels can be detected and decoded using a SDR, a generic active antenna, and a software development kit like gr-iridium [20]. IRIDIUM satellites serve primarily as a communication system. However, the network can also be used for positioning purposes, e.g., through systems like DDK Positioning [21]. IRIDIUM also features a dedicated IRIDIUM IoT service, specifically dedicated to Machine-to-Machine (M2M) applications like transportation, agriculture, oil and gas, utilities, and construction, among many others [22]. In this paper, we use IRIDIUM as the reference LEO satellite constellation to test spoofing attacks because of its popularity, number of users, and availability of an open-source reception module for SDRs to receive real messages, enabling our experimental data collection and analysis.
@@
Note that AEs can also be used as a one-class classification tool, i.e., they can be trained only on samples from one class and then used to detect anomalies as outliers. This is particularly useful when dealing with aerial spoofing attacks. Indeed, collecting attack samples (anomalies) is particularly challenging because of the equipment, technical difficulties required to assemble the experimental setup and launch the attack, and finally, law and safety regulations, which often prevent the collection of such data. Recent related work, such as the one by the authors in [26], also identified experimentally the advantage of autoencoder-based models when a limited number of anomalous samples is available. We use this feature as part of our solution (see Sec. 5).
Paper Location Independence Single Receiver Support for Static Receivers No crypto No Multiple Sources Applied on Satellites No Binary Classification Aerial Adversary
@@
Table 1: Related work on Spoofing Detection.
3 Related Work
@@
Spoofing of satellite messages is usually associated with Global Navigation Satellite System (GNSS) spoofing, i.e., the falsification of positioning and navigation messages emitted by systems such as GPS, GALILEO, and GLONASS [36]. Although secure versions of such communication technologies exist and are used by the military, for backwards-compatibility reasons, civilian systems mostly use their former versions, which are insecure by design. In this context, He et al. in [37] also identified that, similarly to GNSS technologies like GPS, LEO satellite networks typically use a dedicated downlink pilot channel for broadcasting channel status, user management information, call information, etc., as exemplified by Iridium. Attackers can imitate the dedicated pilot channel to broadcast false information to legitimate users, causing network paralysis. However, they did not provide valid countermeasures for such attacks. One common strategy to detect spoofing attacks involves the exploration of alternative sources, such as military GNSS signals [31], terrestrial communication systems [38], or the integration of data from additional GNSS constellations, such as Iridium [18]. These methodologies utilize the correlation between the information collected in distinct channels to assess the authenticity of GNSS signals. However, they depend on the availability and authenticity of an alternative source of information. Some other proposals, such as [30], use multiple receivers to detect spoofing attacks, leveraging spatial and channel diversity and the use of Channel State Information (CSI)-based fingerprints of the received signal. This approach has also been used by the authors in [27] and [32] for WiFi signals, and is overall very promising for indoor spoofing detection. However, as shown in Sec. 7.2, it hardly applies to GNSS spoofing due to high channel noise. To create channel diversity without several receivers, the authors in [33] use a mobile Global Positioning System (GPS) receiver. However, such solutions require mobile receivers and usually incur a non-negligible delay, necessary to actually move from one location to another. Recent popular methods to detect (GNSS) spoofing attacks involve the use of Radio Frequency Fingerprinting (RFF), i.e., the creation of a transmitter fingerprint that helps distinguish legitimate from rogue ones. RFF has been applied successfully for terrestrial communication technologies [39], LEO [9][40][41], and GPS satellites [28]. However, it requires significant training time and does not scale up with the number of legitimate satellites, which is typically large for LEO constellations. Other proposals in the GNSS domain, such as the one by Borhani et al. in [42], use GNSS-specific metrics, such as Cross Ambiguity Function (CAF), to detect spoofing attacks. Being such metrics specific to GNSS technologies, they cannot be used for LEO satellite spoofing detection. Some more recent proposals, such as [10][11], and [29], use IQ samples to detect spoofing attacks affecting both the LEO and GNSS constellations. While the authors in [29] detect anomalies in the Quadrature component of IQ signals, the proposals in [10] and [11] apply binary classification to both I and Q components to detect spoofing and require knowledge of both spoofed and legitimate samples. Binary classification is also required by Kumar et al. in [12], which applied Long Short-Term Memory (LSTM) to noise data via simulations to detect aerial spoofing. It is also worth mentioning the contributions by Abdrabou et al. in [34] and  [35], using Doppler frequency shift (DS) and received power (RP) features to detect spoofing attacks to satellite systems. However, none of these solutions take into account spoofing attacks carried out by aerial attackers, such as drones.
@@
Thus, as summarized in Table 1, the current literature misses actual real-world tests involving spoofing attacks from aerial adversaries. Moreover, existing methods cannot provide a scalable mechanism independent of the number of satellites that comprise the LEO constellation. In Sec. 7.2, we also compare our solution to the solutions proposed in [11] and [9], demonstrating superior performance.
4 System and Adversary Model
@@
Without loss of generality, we consider a satellite transmitter located in the LEO, emitting RF messages according to the IRIDIUM communication system, as discussed in Sec. 2. Thus, legitimate signals emitted by the satellite follow the modulation scheme QPSK with differential encoding, a.k.a. DQPSK, and are transmitted in the frequency band $[1,618.725-1,626.5]$ MHz. The user features a receiving device able to receive and decode IRIDIUM satellite messages, e.g., for communication in hard-to-reach locations. We consider a receiver deployed outdoors for maximum satellite visibility. We do not make any additional assumptions on the deployment of the receiver, which can be either static or mobile. However, we consider that the user does not feature any additional Internet connection than the one possibly provided by the LEO satellite system. Thus, the user cannot resort to any other service than the IRIDIUM communication system to verify the legitimacy of the information provided through such a system. This consideration applies to many IRIDIUM deployments, as the system is typically used to enable communications in remote and hard-to-reach locations. Moreover, note that we require raw physical-layer IQ data to detect spoofing attacks. Such data are usually not available when using COTSs, while they are available by using SDRs, as we did in our experimental campaign. Thus, users willing to deploy our solution require a SDR to collect the data aside to the IRIDIUM receiver provided by the service provider.
@@
5 Satellite Spoofing Detection
In this section, we describe the methodology used to detect spoofing attacks to LEO satellite communication systems. Our proposed methodology for spoofing detection is divided into two phases, i.e., image generation from physical layer information, i.e., IQ samples, (see Sec. 5.1) and spoofing detection (Sec. 5.2). Tab. 2 summarizes our main notation.